OpenStudy (anonymous):

FANS & MEDALS -2(a-8)+6a=36 a=?

OpenStudy (anonymous):

a=5

OpenStudy (anonymous):

and ur correct tjerris

OpenStudy (anonymous):

-2(5-8)+6(5)=36

OpenStudy (anonymous):

we do parenthesis first -2(-3)+30=36

OpenStudy (anonymous):

6+30=36

OpenStudy (anonymous):

Btw thanks for the medal myronnapora

OpenStudy (anonymous):

np wnt another one

OpenStudy (anonymous):

perenthisi first and 5-8=-3 then -2 times -3 equals 6 then multiply and 6 time 5 = 30 then add 30 + 6 =36

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yes u got it on the spot